OpenStudy (anonymous):

I need some extra brainpower for this: a music site observed that 60% of the people who downloaded songs from its site downloaded Q Sam's latest single. The equation below represents this information, where x represents the total number of people who downloaded songs from the site that month: x = 0.6x + 384 (WHAT IVE WORKED OUT): X = 100% Of people who downloaded songs from a website 0.6 [times] X = 60% 60% + 384 = 100% Choices: How many people who downloaded songs from the site that month downloaded Q Sam's latest single? A. 576 B. 240 C. 144 D. 960

OpenStudy (amistre64):

x = 0.6x + 384 -.6x -.6x ----------------- .4x = 384 what should x be?

OpenStudy (anonymous):

96

OpenStudy (amistre64):

almost .. you forgot a 0

OpenStudy (amistre64):

384/.4 = 960

OpenStudy (anonymous):

!!!!

OpenStudy (anonymous):

You mind

OpenStudy (anonymous):

helping out on some others

OpenStudy (amistre64):

dunno, i havent seen the others ... they might scare me

OpenStudy (anonymous):

heres one

OpenStudy (anonymous):

Solve for : x + 20 / 2 = 3x

OpenStudy (amistre64):

use your grouping symbols correctly ... im assuming (x+20) is on top?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

what is your thoughts then, how would you try to work this?

OpenStudy (anonymous):

Ill show you

OpenStudy (anonymous):

Wait... Ima do this on notepad

OpenStudy (anonymous):

I know that you first have to single x out onto 1 side of the line, but what confuses me is the division thing. I know with division you have to find a Least common denominator, but since there is only one division problem here, Im clueless as how to go on.

OpenStudy (amistre64):

whatever we do to one side, we do to the other .... A/2 = B ... in order to UNDO division, we multiply 2A/2 = 2B A = 2B

OpenStudy (amistre64):

i would start out by multiplying both sides by that denominator

OpenStudy (anonymous):

So multiply both sides by 2 and get rid of the 2 on the first side?

OpenStudy (amistre64):

yep x+20 = 2(3x) x+20 = 6x ... now what are your thoughts?

OpenStudy (anonymous):

There is an x on both sides, so you have to get them onto one side. Which side do I choose?

OpenStudy (amistre64):

either side works ... personally is move the leftside over but it doesnt matter

OpenStudy (anonymous):

20 = 5x?

OpenStudy (amistre64):

subtraction UNDOES addition soo x+20 = 6x -x -x ----------- 20 = 5x yes, now what is our x value?

OpenStudy (anonymous):

4

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

woww

OpenStudy (amistre64):

if we had options, we could have also tried them to see what worked

OpenStudy (anonymous):

I have one more. You wanna help pal?

OpenStudy (amistre64):

one more, fine

OpenStudy (anonymous):

Here I will snip and send pic

OpenStudy (amistre64):

i few things wrong with this ... some would contend that this is cheating on an exam; but the more important thing is that you have posted personal information that is contained in the picture.

OpenStudy (amistre64):

anyways, you know how to distribute?

OpenStudy (anonymous):

Its no exam, XD Its a module test.

OpenStudy (anonymous):

3x + 3 = -2x - 2 + 6

OpenStudy (amistre64):

mrs collins, santiago, seminole county ... none of that needs to be posted publically is all. its a safety issue for minors

OpenStudy (amistre64):

-2(-1) is not -2

OpenStudy (amistre64):

dont forget your negative rules :)

OpenStudy (anonymous):

Seriously? I could give out my homestate and city, there would still be about 1000+ collins and Santiagos to pick fro. That and you dont even know my address.

OpenStudy (amistre64):

3x + 3 = -2x + 2 + 6 now what?

OpenStudy (anonymous):

ANYWAYS

OpenStudy (anonymous):

3x + 3 = -2x + 2 + 6 3x + 3 = -2x + 8 3x = -2x + 5?

OpenStudy (amistre64):

so far so good, now add 2x to get rid of the -2x and we should be good

OpenStudy (anonymous):

3(x + 1) = −2(x − 1) + 6. 3x + 3 = -2x + 2 + 6 3x + 3 = -2x + 8 3x = -2x + 5? +2x +2x 5x = 5 x = 1 Thanks pal! :D

OpenStudy (amistre64):

good luck :)